OpenStudy (anonymous):

if half the rangeof the increase series11,x,23,y,z,68,73 is equal to its median what is the median of the series?

OpenStudy (anonymous):

help me in my question

OpenStudy (anonymous):

range = highest obs - lowest obs = 73 - 11 = 62 half of range = 62/2 = 31 so median is 31

OpenStudy (anonymous):

Hence y = 31 x >= 11 and x <=23 and z >= 31 and <= 68

OpenStudy (anonymous):

if n is an integer between 0 and 100 then any of the following could be 3n+3 expect?

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

I don't really understand the question, also if you post it as a new question then more people will see it